OpenStudy (vianne):

Which literal equations are equivalent to m = W/g? Choose all answers that are correct. A. g = W/m B. g = m/W C. W = mg D. W = m/g

OpenStudy (anonymous):

you have replied enough there and now come back. :P

OpenStudy (anonymous):

\[m = \frac{W}{g} \]

OpenStudy (vianne):

:33 Was having a little break. ;)

OpenStudy (vianne):

Anyways, I don't get how to solve this problem?

OpenStudy (anonymous):

Don't let this break become "Permanent".. :P \(\dagger\).

OpenStudy (anonymous):

In your answer options, either you have \(g\) or you have \(W\) in left hand side? Right?

OpenStudy (vianne):

:P Getting spammed with notifications. ;-;

OpenStudy (vianne):

I thought it was right.. Since.. m is on the left side of the equal sign..

OpenStudy (anonymous):

So, we do it first for \(g\).

OpenStudy (vianne):

k

OpenStudy (anonymous):

you have : \[m= \frac{W}{g}\] If you will multiply by \(g\) on both the sides, then what will you get?

OpenStudy (vianne):

mg and wg right?

OpenStudy (anonymous):

\(mg\) on left is right but think it again do you really get \(wg\) on right?

OpenStudy (vianne):

Erm.. No, because you have W dividing g?

OpenStudy (anonymous):

What about the other \(g\) which you have in denominator? why you don't like that \(g\) ?

OpenStudy (anonymous):

When you will multiply by \(g\) on both sides, you will have something like this: \[m \color{green}{\times g} = \frac{W}{g} \color{green}{\times g}\]

OpenStudy (anonymous):

Right?

OpenStudy (vianne):

right

OpenStudy (anonymous):

Don't you think \(g\) and \(g\) will cancel out?

OpenStudy (anonymous):

\[m g = \frac{W}{\cancel{g}} \times \cancel{g}\]

OpenStudy (vianne):

Was going to write a stupid reply. :P "H is going to be next for g x g"

OpenStudy (anonymous):

Hooray..!! I was doing it for \(g\) but we have luckily found \(W\). So, \(W = mg\) agree?

OpenStudy (vianne):

Yep

OpenStudy (anonymous):

Sorry, look at your answer option again, you have written something wrong there.

OpenStudy (anonymous):

Both C and D options are one and the same, check that again.

OpenStudy (vianne):

heh... C. W = mg ;3

OpenStudy (vianne):

I accidently typed it wrong..

OpenStudy (anonymous):

So, one correct option we have found, okay?

OpenStudy (vianne):

k

OpenStudy (vianne):

This keyboard sucks..

OpenStudy (anonymous):

Now you have : \(mg = W\) Now, tell me what will you get when you divide by \(m\) both the sides?

OpenStudy (vianne):

You'd get W/m and m/g

OpenStudy (anonymous):

Then what are you waiting for, take them and enjoy.. :P \(\dagger\).. :P

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

m/g ??

OpenStudy (vianne):

um..

OpenStudy (anonymous):

How about breaking that keyboard? :P

OpenStudy (vianne):

I'm using my older brother's keyboard..

OpenStudy (vianne):

So it'd be A. g = W/m right?

OpenStudy (anonymous):

Right.

OpenStudy (vianne):

because if you do m divide m = crossed out so you'd only have g left.

OpenStudy (anonymous):

And ?

OpenStudy (anonymous):

A and??

OpenStudy (vianne):

And and and C

OpenStudy (anonymous):

make sure C is W = mg. and not W = m/g. Okay?

OpenStudy (vianne):

Okay, I editted the question so it'd look like W = mg ;0

OpenStudy (vianne):

Next Question Shall Be On Another Post... Because.. I want to double check something.. With you.

OpenStudy (anonymous):

Sure, you can triple check also, if you want. :P